Find the whole number such that twicw its square is 11 more than 21 times the original number ?
Answer by nerdybill(7384)

Let x = the whole number
.
From: "twice its square is 11 more than 21 times the original number"
2x^2 = 21x+11
2x^2 - 21x - 11 = 0
Factoring the left:
(2x+1)(x-11) = 0
.
Setting each term on the left to zero:
(2x+1) = 0
2x = -1
x = -1/2
.
(x-11) = 0
x = 11
.
Solution set:
x = {-1/2, 11}
.
Since the problem asks for a "whole number" we eliminate the -1/2 leaving us with:
x = 11